Question #211163

Find the probability that a randomly selected senior high school student spends less than 21 hours or greater than 30 hours


Expert's answer

Mean is 24 hours with a standard deviation of 4 hours.


Let's find the opposite (the probability that a student spends more than 21 and less than 30 hours), than subtract this probability from 1, and that will be the answer.


The probability that 21 < X < 30 is equal to the blue area under the curve.





Since μ\mu = 24 and σ=4\sigma =4  we have:

P( 21 < X < 30 ) = P(21-24 < X - μ\mu < 30 - 24) = P(21−244<X−μσ<30−244)(\frac{21-24}{4}<\frac{X-\mu}{\sigma} <\frac{30-24}{4} )

Since Z=X−μσZ = \frac{X-\mu}{\sigma}, 21−244=−0.75\frac{21-24}{4} = -0.75 and 30−244=1.5\frac{30-24}{4} = 1.5 we have:

P( 21 < X < 30) = P(-0.75 < Z < 1.5)

Use the standard normal table to conclude that:

P(-0.75 < Z < 1.5) = 0.7066

So, P' = 1 - 0.7066 = 0.2934

Answer: P = 0.2934.


P.S. Here is the z-score table:



LATEST TUTORIALS
APPROVED BY CLIENTS